\providecommand{\href}[2]{#2}\begingroup\raggedright\begin{thebibliography}{10}

\bibitem{LHCReinterpretationForum:2020xtr}
{\scshape LHC Reinterpretation Forum} collaboration, W.~Abdallah et~al.,
  \emph{{Reinterpretation of LHC Results for New Physics: Status and
  Recommendations after Run 2}},
  \href{https://doi.org/10.21468/SciPostPhys.9.2.022}{\emph{SciPost Phys.}
  {\bfseries 9} (2020) 022},
  [\href{https://arxiv.org/abs/2003.07868}{{\ttfamily 2003.07868}}].

\bibitem{Kraml:2013mwa}
S.~Kraml, S.~Kulkarni, U.~Laa, A.~Lessa, W.~Magerl, D.~Proschofsky-Spindler
  et~al., \emph{{SModelS: a tool for interpreting simplified-model results from
  the LHC and its application to supersymmetry}},
  \href{https://doi.org/10.1140/epjc/s10052-014-2868-5}{\emph{Eur. Phys. J. C}
  {\bfseries 74} (2014) 2868},
  [\href{https://arxiv.org/abs/1312.4175}{{\ttfamily 1312.4175}}].

\bibitem{Alguero:2020grj}
G.~Alguero, S.~Kraml and W.~Waltenberger, \emph{{A SModelS interface for pyhf
  likelihoods}}, \href{https://doi.org/10.1016/j.cpc.2021.107909}{\emph{Comput.
  Phys. Commun.} {\bfseries 264} (2021) 107909},
  [\href{https://arxiv.org/abs/2009.01809}{{\ttfamily 2009.01809}}].

\bibitem{Ilten:2018crw}
P.~Ilten, Y.~Soreq, M.~Williams and W.~Xue, \emph{{Serendipity in dark photon
  searches}}, \href{https://doi.org/10.1007/JHEP06(2018)004}{\emph{JHEP}
  {\bfseries 06} (2018) 004},
  [\href{https://arxiv.org/abs/1801.04847}{{\ttfamily 1801.04847}}].

\bibitem{Sjostrand:2007gs}
T.~Sjostrand, S.~Mrenna and P.~Z. Skands, \emph{{A Brief Introduction to PYTHIA
  8.1}}, \href{https://doi.org/10.1016/j.cpc.2008.01.036}{\emph{Comput. Phys.
  Commun.} {\bfseries 178} (2008) 852--867},
  [\href{https://arxiv.org/abs/0710.3820}{{\ttfamily 0710.3820}}].

\bibitem{deFavereau:2013fsa}
{\scshape DELPHES 3} collaboration, J.~de~Favereau, C.~Delaere, P.~Demin,
  A.~Giammanco, V.~Lema\^\i{}tre, A.~Mertens et~al., \emph{{DELPHES 3, A
  modular framework for fast simulation of a generic collider experiment}},
  \href{https://doi.org/10.1007/JHEP02(2014)057}{\emph{JHEP} {\bfseries 02}
  (2014) 057}, [\href{https://arxiv.org/abs/1307.6346}{{\ttfamily 1307.6346}}].

\bibitem{Conte:2012fm}
E.~Conte, B.~Fuks and G.~Serret, \emph{{MadAnalysis 5, A User-Friendly
  Framework for Collider Phenomenology}},
  \href{https://doi.org/10.1016/j.cpc.2012.09.009}{\emph{Comput. Phys. Commun.}
  {\bfseries 184} (2013) 222--256},
  [\href{https://arxiv.org/abs/1206.1599}{{\ttfamily 1206.1599}}].

\bibitem{Conte:2014zja}
E.~Conte, B.~Dumont, B.~Fuks and C.~Wymant, \emph{{Designing and recasting LHC
  analyses with MadAnalysis 5}},
  \href{https://doi.org/10.1140/epjc/s10052-014-3103-0}{\emph{Eur. Phys. J. C}
  {\bfseries 74} (2014) 3103},
  [\href{https://arxiv.org/abs/1405.3982}{{\ttfamily 1405.3982}}].

\bibitem{Dumont:2014tja}
B.~Dumont, B.~Fuks, S.~Kraml, S.~Bein, G.~Chalons, E.~Conte et~al.,
  \emph{{Toward a public analysis database for LHC new physics searches using
  MADANALYSIS 5}},
  \href{https://doi.org/10.1140/epjc/s10052-014-3242-3}{\emph{Eur. Phys. J. C}
  {\bfseries 75} (2015) 56}, [\href{https://arxiv.org/abs/1407.3278}{{\ttfamily
  1407.3278}}].

\bibitem{Conte:2018vmg}
E.~Conte and B.~Fuks, \emph{{Confronting new physics theories to LHC data with
  MADANALYSIS 5}}, \href{https://doi.org/10.1142/S0217751X18300272}{\emph{Int.
  J. Mod. Phys. A} {\bfseries 33} (2018) 1830027},
  [\href{https://arxiv.org/abs/1808.00480}{{\ttfamily 1808.00480}}].

\bibitem{Araz:2019otb}
J.~Y. Araz, M.~Frank and B.~Fuks, \emph{{Reinterpreting the results of the LHC
  with MadAnalysis 5: uncertainties and higher-luminosity estimates}},
  \href{https://doi.org/10.1140/epjc/s10052-020-8076-6}{\emph{Eur. Phys. J. C}
  {\bfseries 80} (2020) 531},
  [\href{https://arxiv.org/abs/1910.11418}{{\ttfamily 1910.11418}}].

\bibitem{Araz:2020lnp}
J.~Y. Araz, B.~Fuks and G.~Polykratis, \emph{{Simplified fast detector
  simulation in MADANALYSIS 5}},
  \href{https://doi.org/10.1140/epjc/s10052-021-09052-5}{\emph{Eur. Phys. J. C}
  {\bfseries 81} (2021) 329},
  [\href{https://arxiv.org/abs/2006.09387}{{\ttfamily 2006.09387}}].

\bibitem{Drees:2013wra}
M.~Drees, H.~Dreiner, D.~Schmeier, J.~Tattersall and J.~S. Kim,
  \emph{{CheckMATE: Confronting your Favourite New Physics Model with LHC
  Data}}, \href{https://doi.org/10.1016/j.cpc.2014.10.018}{\emph{Comput. Phys.
  Commun.} {\bfseries 187} (2015) 227--265},
  [\href{https://arxiv.org/abs/1312.2591}{{\ttfamily 1312.2591}}].

\bibitem{Dercks:2016npn}
D.~Dercks, N.~Desai, J.~S. Kim, K.~Rolbiecki, J.~Tattersall and T.~Weber,
  \emph{{CheckMATE 2: From the model to the limit}},
  \href{https://doi.org/10.1016/j.cpc.2017.08.021}{\emph{Comput. Phys. Commun.}
  {\bfseries 221} (2017) 383--418},
  [\href{https://arxiv.org/abs/1611.09856}{{\ttfamily 1611.09856}}].

\bibitem{GAMBIT:2017qxg}
{\scshape GAMBIT} collaboration, C.~Bal\'azs et~al., \emph{{ColliderBit: a
  GAMBIT module for the calculation of high-energy collider observables and
  likelihoods}},
  \href{https://doi.org/10.1140/epjc/s10052-017-5285-8}{\emph{Eur. Phys. J. C}
  {\bfseries 77} (2017) 795},
  [\href{https://arxiv.org/abs/1705.07919}{{\ttfamily 1705.07919}}].

\bibitem{GAMBIT:2017yxo}
{\scshape GAMBIT} collaboration, P.~Athron et~al., \emph{{GAMBIT: The Global
  and Modular Beyond-the-Standard-Model Inference Tool}},
  \href{https://doi.org/10.1140/epjc/s10052-017-5321-8}{\emph{Eur. Phys. J. C}
  {\bfseries 77} (2017) 784},
  [\href{https://arxiv.org/abs/1705.07908}{{\ttfamily 1705.07908}}].

\bibitem{Buckley:2010ar}
A.~Buckley, J.~Butterworth, D.~Grellscheid, H.~Hoeth, L.~Lonnblad, J.~Monk
  et~al., \emph{{Rivet user manual}},
  \href{https://doi.org/10.1016/j.cpc.2013.05.021}{\emph{Comput. Phys. Commun.}
  {\bfseries 184} (2013) 2803--2819},
  [\href{https://arxiv.org/abs/1003.0694}{{\ttfamily 1003.0694}}].

\bibitem{Bierlich:2019rhm}
C.~Bierlich et~al., \emph{{Robust Independent Validation of Experiment and
  Theory: Rivet version 3}},
  \href{https://doi.org/10.21468/SciPostPhys.8.2.026}{\emph{SciPost Phys.}
  {\bfseries 8} (2020) 026},
  [\href{https://arxiv.org/abs/1912.05451}{{\ttfamily 1912.05451}}].

\bibitem{Butterworth:2016sqg}
J.~M. Butterworth, D.~Grellscheid, M.~Kr\"amer, B.~Sarrazin and D.~Yallup,
  \emph{{Constraining new physics with collider measurements of Standard Model
  signatures}}, \href{https://doi.org/10.1007/JHEP03(2017)078}{\emph{JHEP}
  {\bfseries 03} (2017) 078},
  [\href{https://arxiv.org/abs/1606.05296}{{\ttfamily 1606.05296}}].

\bibitem{Amrith:2018yfb}
S.~Amrith, J.~M. Butterworth, F.~F. Deppisch, W.~Liu, A.~Varma and D.~Yallup,
  \emph{{LHC Constraints on a $B-L$ Gauge Model using Contur}},
  \href{https://doi.org/10.1007/JHEP05(2019)154}{\emph{JHEP} {\bfseries 05}
  (2019) 154}, [\href{https://arxiv.org/abs/1811.11452}{{\ttfamily
  1811.11452}}].

\bibitem{Buckley:2020wzk}
A.~Buckley, J.~M. Butterworth, L.~Corpe, D.~Huang and P.~Sun, \emph{{New
  sensitivity of current LHC measurements to vector-like quarks}},
  \href{https://doi.org/10.21468/SciPostPhys.9.5.069}{\emph{SciPost Phys.}
  {\bfseries 9} (2020) 069},
  [\href{https://arxiv.org/abs/2006.07172}{{\ttfamily 2006.07172}}].

\bibitem{Butterworth:2020vnb}
J.~M. Butterworth, M.~Habedank, P.~Pani and A.~Vaitkus, \emph{{A study of
  collider signatures for two Higgs doublet models with a Pseudoscalar mediator
  to Dark Matter}},
  \href{https://doi.org/10.21468/SciPostPhysCore.4.1.003}{\emph{SciPost Phys.
  Core} {\bfseries 4} (2021) 003},
  [\href{https://arxiv.org/abs/2009.02220}{{\ttfamily 2009.02220}}].

\bibitem{Buckley:2021neu}
A.~Buckley et~al., \emph{{Testing new physics models with global comparisons to
  collider measurements: the Contur toolkit}},
  \href{https://doi.org/10.21468/SciPostPhysCore.4.2.013}{\emph{SciPost Phys.
  Core} {\bfseries 4} (2021) 013},
  [\href{https://arxiv.org/abs/2102.04377}{{\ttfamily 2102.04377}}].

\bibitem{CMS:2018kag}
{\scshape CMS} collaboration, A.~M. Sirunyan et~al., \emph{{Search for new
  physics in events with two soft oppositely charged leptons and missing
  transverse momentum in proton-proton collisions at $\sqrt{s}=$ 13 TeV}},
  \href{https://doi.org/10.1016/j.physletb.2018.05.062}{\emph{Phys. Lett. B}
  {\bfseries 782} (2018) 440--467},
  [\href{https://arxiv.org/abs/1801.01846}{{\ttfamily 1801.01846}}].

\bibitem{Brooijmans:2020yij}
G.~Brooijmans et~al., \emph{{Les Houches 2019 Physics at TeV Colliders: New
  Physics Working Group Report}},  in \emph{{11th Les Houches Workshop on
  Physics at TeV Colliders}: {PhysTeV Les Houches}}, 2, 2020,
  \href{https://arxiv.org/abs/2002.12220}{{\ttfamily 2002.12220}}.

\bibitem{Arina:2020alg}
C.~Arina, \emph{{Review on Dark Matter Tools}},
  \href{https://doi.org/10.22323/1.392.0001}{\emph{PoS} {\bfseries TOOLS2020}
  (2021) 001}, [\href{https://arxiv.org/abs/2012.09462}{{\ttfamily
  2012.09462}}].

\bibitem{Gondolo:2004sc}
P.~Gondolo, J.~Edsjo, P.~Ullio, L.~Bergstrom, M.~Schelke and E.~A. Baltz,
  \emph{{DarkSUSY: Computing supersymmetric dark matter properties
  numerically}},
  \href{https://doi.org/10.1088/1475-7516/2004/07/008}{\emph{JCAP} {\bfseries
  07} (2004) 008}, [\href{https://arxiv.org/abs/astro-ph/0406204}{{\ttfamily
  astro-ph/0406204}}].

\bibitem{Bringmann:2018lay}
T.~Bringmann, J.~Edsj\"o, P.~Gondolo, P.~Ullio and L.~Bergstr\"om,
  \emph{{DarkSUSY 6 : An Advanced Tool to Compute Dark Matter Properties
  Numerically}},
  \href{https://doi.org/10.1088/1475-7516/2018/07/033}{\emph{JCAP} {\bfseries
  07} (2018) 033}, [\href{https://arxiv.org/abs/1802.03399}{{\ttfamily
  1802.03399}}].

\bibitem{Belanger:2001fz}
G.~Belanger, F.~Boudjema, A.~Pukhov and A.~Semenov, \emph{{MicrOMEGAs: A
  Program for calculating the relic density in the MSSM}},
  \href{https://doi.org/10.1016/S0010-4655(02)00596-9}{\emph{Comput. Phys.
  Commun.} {\bfseries 149} (2002) 103--120},
  [\href{https://arxiv.org/abs/hep-ph/0112278}{{\ttfamily hep-ph/0112278}}].

\bibitem{Belanger:2004yn}
G.~Belanger, F.~Boudjema, A.~Pukhov and A.~Semenov, \emph{{micrOMEGAs: Version
  1.3}}, \href{https://doi.org/10.1016/j.cpc.2005.12.005}{\emph{Comput. Phys.
  Commun.} {\bfseries 174} (2006) 577--604},
  [\href{https://arxiv.org/abs/hep-ph/0405253}{{\ttfamily hep-ph/0405253}}].

\bibitem{Belanger:2006is}
G.~Belanger, F.~Boudjema, A.~Pukhov and A.~Semenov, \emph{{MicrOMEGAs 2.0: A
  Program to calculate the relic density of dark matter in a generic model}},
  \href{https://doi.org/10.1016/j.cpc.2006.11.008}{\emph{Comput. Phys. Commun.}
  {\bfseries 176} (2007) 367--382},
  [\href{https://arxiv.org/abs/hep-ph/0607059}{{\ttfamily hep-ph/0607059}}].

\bibitem{Belanger:2008sj}
G.~Belanger, F.~Boudjema, A.~Pukhov and A.~Semenov, \emph{{Dark matter direct
  detection rate in a generic model with micrOMEGAs 2.2}},
  \href{https://doi.org/10.1016/j.cpc.2008.11.019}{\emph{Comput. Phys. Commun.}
  {\bfseries 180} (2009) 747--767},
  [\href{https://arxiv.org/abs/0803.2360}{{\ttfamily 0803.2360}}].

\bibitem{Belanger:2010gh}
G.~Belanger, F.~Boudjema, P.~Brun, A.~Pukhov, S.~Rosier-Lees, P.~Salati et~al.,
  \emph{{Indirect search for dark matter with micrOMEGAs2.4}},
  \href{https://doi.org/10.1016/j.cpc.2010.11.033}{\emph{Comput. Phys. Commun.}
  {\bfseries 182} (2011) 842--856},
  [\href{https://arxiv.org/abs/1004.1092}{{\ttfamily 1004.1092}}].

\bibitem{Belanger:2013oya}
G.~Belanger, F.~Boudjema, A.~Pukhov and A.~Semenov, \emph{{micrOMEGAs$\_$3: A
  program for calculating dark matter observables}},
  \href{https://doi.org/10.1016/j.cpc.2013.10.016}{\emph{Comput. Phys. Commun.}
  {\bfseries 185} (2014) 960--985},
  [\href{https://arxiv.org/abs/1305.0237}{{\ttfamily 1305.0237}}].

\bibitem{Belanger:2014vza}
G.~B\'elanger, F.~Boudjema, A.~Pukhov and A.~Semenov, \emph{{micrOMEGAs4.1: two
  dark matter candidates}},
  \href{https://doi.org/10.1016/j.cpc.2015.03.003}{\emph{Comput. Phys. Commun.}
  {\bfseries 192} (2015) 322--329},
  [\href{https://arxiv.org/abs/1407.6129}{{\ttfamily 1407.6129}}].

\bibitem{Backovic:2013dpa}
M.~Backovic, K.~Kong and M.~McCaskey, \emph{{MadDM v.1.0: Computation of Dark
  Matter Relic Abundance Using MadGraph5}},
  \href{https://doi.org/10.1016/j.dark.2014.04.001}{\emph{Physics of the Dark
  Universe} {\bfseries 5-6} (2014) 18--28},
  [\href{https://arxiv.org/abs/1308.4955}{{\ttfamily 1308.4955}}].

\bibitem{Backovic:2015cra}
M.~Backovi\'c, A.~Martini, O.~Mattelaer, K.~Kong and G.~Mohlabeng,
  \emph{{Direct Detection of Dark Matter with MadDM v.2.0}},
  \href{https://doi.org/10.1016/j.dark.2015.09.001}{\emph{Phys. Dark Univ.}
  {\bfseries 9-10} (5, 2015) 37--50},
  [\href{https://arxiv.org/abs/1505.04190}{{\ttfamily 1505.04190}}].

\bibitem{Ambrogi:2018jqj}
F.~Ambrogi, C.~Arina, M.~Backovic, J.~Heisig, F.~Maltoni, L.~Mantani et~al.,
  \emph{{MadDM v.3.0: a Comprehensive Tool for Dark Matter Studies}},
  \href{https://doi.org/10.1016/j.dark.2018.11.009}{\emph{Phys. Dark Univ.}
  {\bfseries 24} (2019) 100249},
  [\href{https://arxiv.org/abs/1804.00044}{{\ttfamily 1804.00044}}].

\bibitem{Arina:2020kko}
C.~Arina, J.~Heisig, F.~Maltoni, L.~Mantani, D.~Massaro, O.~Mattelaer et~al.,
  \emph{{Studying dark matter with MadDM 3.1: a short user guide}},
  \href{https://doi.org/10.22323/1.392.0009}{\emph{PoS} {\bfseries TOOLS2020}
  (2021) 009}, [\href{https://arxiv.org/abs/2012.09016}{{\ttfamily
  2012.09016}}].

\bibitem{Arbey:2009gu}
A.~Arbey and F.~Mahmoudi, \emph{{SuperIso Relic: A Program for calculating
  relic density and flavor physics observables in Supersymmetry}},
  \href{https://doi.org/10.1016/j.cpc.2010.03.010}{\emph{Comput. Phys. Commun.}
  {\bfseries 181} (2010) 1277--1292},
  [\href{https://arxiv.org/abs/0906.0369}{{\ttfamily 0906.0369}}].

\bibitem{Arbey:2011zz}
A.~Arbey and F.~Mahmoudi, \emph{{SuperIso Relic v3.0: A program for calculating
  relic density and flavour physics observables: Extension to NMSSM}},
  \href{https://doi.org/10.1016/j.cpc.2011.03.019}{\emph{Comput. Phys. Commun.}
  {\bfseries 182} (2011) 1582--1583}.

\bibitem{Arbey:2018msw}
A.~Arbey, F.~Mahmoudi and G.~Robbins, \emph{{SuperIso Relic v4: A program for
  calculating dark matter and flavour physics observables in Supersymmetry}},
  \href{https://doi.org/10.1016/j.cpc.2019.01.014}{\emph{Comput. Phys. Commun.}
  {\bfseries 239} (2019) 238--264},
  [\href{https://arxiv.org/abs/1806.11489}{{\ttfamily 1806.11489}}].

\bibitem{GAMBITDarkMatterWorkgroup:2017fax}
{\scshape GAMBIT Dark Matter Workgroup} collaboration, T.~Bringmann et~al.,
  \emph{{DarkBit: A GAMBIT module for computing dark matter observables and
  likelihoods}},
  \href{https://doi.org/10.1140/epjc/s10052-017-5155-4}{\emph{Eur. Phys. J. C}
  {\bfseries 77} (2017) 831},
  [\href{https://arxiv.org/abs/1705.07920}{{\ttfamily 1705.07920}}].

\bibitem{GAMBIT:2018eea}
{\scshape GAMBIT} collaboration, P.~Athron et~al., \emph{{Global analyses of
  Higgs portal singlet dark matter models using GAMBIT}},
  \href{https://doi.org/10.1140/epjc/s10052-018-6513-6}{\emph{Eur. Phys. J. C}
  {\bfseries 79} (2019) 38},
  [\href{https://arxiv.org/abs/1808.10465}{{\ttfamily 1808.10465}}].

\bibitem{Bishara:2017nnn}
F.~Bishara, J.~Brod, B.~Grinstein and J.~Zupan, \emph{{DirectDM: a tool for
  dark matter direct detection}},
  \href{https://arxiv.org/abs/1708.02678}{{\ttfamily 1708.02678}}.

\bibitem{Brod:2017bsw}
J.~Brod, A.~Gootjes-Dreesbach, M.~Tammaro and J.~Zupan, \emph{{Effective Field
  Theory for Dark Matter Direct Detection up to Dimension Seven}},
  \href{https://doi.org/10.1007/JHEP10(2018)065}{\emph{JHEP} {\bfseries 10}
  (2018) 065}, [\href{https://arxiv.org/abs/1710.10218}{{\ttfamily
  1710.10218}}].

\bibitem{Kozar:2021iur}
N.~A. Kozar, A.~Caddell, L.~Fraser-Leach, P.~Scott and A.~C. Vincent,
  \emph{{Capt'n General: A generalized stellar dark matter capture and heat
  transport code}},  in \emph{{Tools for High Energy Physics and Cosmology}},
  5, 2021, \href{https://arxiv.org/abs/2105.06810}{{\ttfamily 2105.06810}}.

\bibitem{IceCube:2012fvn}
{\scshape IceCube} collaboration, P.~Scott et~al., \emph{{Use of event-level
  neutrino telescope data in global fits for theories of new physics}},
  \href{https://doi.org/10.1088/1475-7516/2012/11/057}{\emph{JCAP} {\bfseries
  11} (2012) 057}, [\href{https://arxiv.org/abs/1207.0810}{{\ttfamily
  1207.0810}}].

\bibitem{IceCube:2016yoy}
{\scshape IceCube} collaboration, M.~G. Aartsen et~al., \emph{{Improved limits
  on dark matter annihilation in the Sun with the 79-string IceCube detector
  and implications for supersymmetry}},
  \href{https://doi.org/10.1088/1475-7516/2016/04/022}{\emph{JCAP} {\bfseries
  04} (2016) 022}, [\href{https://arxiv.org/abs/1601.00653}{{\ttfamily
  1601.00653}}].

\bibitem{Stocker:2018avm}
P.~St\"ocker, M.~Kr\"amer, J.~Lesgourgues and V.~Poulin, \emph{{Exotic energy
  injection with ExoCLASS: Application to the Higgs portal model and
  evaporating black holes}},
  \href{https://doi.org/10.1088/1475-7516/2018/03/018}{\emph{JCAP} {\bfseries
  03} (2018) 018}, [\href{https://arxiv.org/abs/1801.01871}{{\ttfamily
  1801.01871}}].

\bibitem{Blas:2011aaa}
D.~{Blas}, J.~{Lesgourgues} and T.~{Tram}, \emph{{The Cosmic Linear Anisotropy
  Solving System (CLASS). Part II: Approximation schemes}},
  \href{https://doi.org/10.1088/1475-7516/2011/07/034}{\emph{JCAP} {\bfseries
  2011} (July, 2011) 034}, [\href{https://arxiv.org/abs/1104.2933}{{\ttfamily
  1104.2933}}].

\bibitem{DarkMachinesHighDimensionalSamplingGroup:2021wkt}
{\scshape DarkMachines High Dimensional Sampling Group} collaboration,
  C.~Bal\'azs et~al., \emph{{A comparison of optimisation algorithms for
  high-dimensional particle and astrophysics applications}},
  \href{https://doi.org/10.1007/JHEP05(2021)108}{\emph{JHEP} {\bfseries 05}
  (2021) 108}, [\href{https://arxiv.org/abs/2101.04525}{{\ttfamily
  2101.04525}}].

\bibitem{AbdusSalam:2020rdj}
S.~S. AbdusSalam et~al., \emph{{Simple and statistically sound strategies for
  analysing physical theories}},
  \href{https://arxiv.org/abs/2012.09874}{{\ttfamily 2012.09874}}.

\bibitem{Bechtle:2012zk}
P.~Bechtle et~al., \emph{{Constrained Supersymmetry after two years of LHC
  data: a global view with Fittino}},
  \href{https://doi.org/10.1007/JHEP06(2012)098}{\emph{JHEP} {\bfseries 06}
  (2012) 098}, [\href{https://arxiv.org/abs/1204.4199}{{\ttfamily 1204.4199}}].

\bibitem{RuizdeAustri:2006iwb}
R.~Ruiz~de Austri, R.~Trotta and L.~Roszkowski, \emph{{A Markov chain Monte
  Carlo analysis of the CMSSM}},
  \href{https://doi.org/10.1088/1126-6708/2006/05/002}{\emph{JHEP} {\bfseries
  05} (2006) 002}, [\href{https://arxiv.org/abs/hep-ph/0602028}{{\ttfamily
  hep-ph/0602028}}].

\bibitem{Han:2016gvr}
C.~Han, K.-i. Hikasa, L.~Wu, J.~M. Yang and Y.~Zhang, \emph{{Status of CMSSM in
  light of current LHC Run-2 and LUX data}},
  \href{https://doi.org/10.1016/j.physletb.2017.04.026}{\emph{Phys. Lett. B}
  {\bfseries 769} (2017) 470--476},
  [\href{https://arxiv.org/abs/1612.02296}{{\ttfamily 1612.02296}}].

\bibitem{DeBlas:2019ehy}
J.~De~Blas et~al., \emph{{$\texttt{HEPfit}$: a code for the combination of
  indirect and direct constraints on high energy physics models}},
  \href{https://doi.org/10.1140/epjc/s10052-020-7904-z}{\emph{Eur. Phys. J. C}
  {\bfseries 80} (2020) 456},
  [\href{https://arxiv.org/abs/1910.14012}{{\ttfamily 1910.14012}}].

\bibitem{Chiang:2018cgb}
C.-W. Chiang, G.~Cottin and O.~Eberhardt, \emph{{Global fits in the
  Georgi-Machacek model}},
  \href{https://doi.org/10.1103/PhysRevD.99.015001}{\emph{Phys. Rev. D}
  {\bfseries 99} (2019) 015001},
  [\href{https://arxiv.org/abs/1807.10660}{{\ttfamily 1807.10660}}].

\bibitem{Chowdhury:2017aav}
D.~Chowdhury and O.~Eberhardt, \emph{{Update of Global Two-Higgs-Doublet Model
  Fits}}, \href{https://doi.org/10.1007/JHEP05(2018)161}{\emph{JHEP} {\bfseries
  05} (2018) 161}, [\href{https://arxiv.org/abs/1711.02095}{{\ttfamily
  1711.02095}}].

\bibitem{Kvellestad:2019vxm}
A.~Kvellestad, P.~Scott and M.~White, \emph{{GAMBIT and its Application in the
  Search for Physics Beyond the Standard Model}},
  \href{https://arxiv.org/abs/1912.04079}{{\ttfamily 1912.04079}}.

\bibitem{GAMBITFlavourWorkgroup:2017dbx}
{\scshape GAMBIT Flavour Workgroup} collaboration, F.~U. Bernlochner et~al.,
  \emph{{FlavBit: A GAMBIT module for computing flavour observables and
  likelihoods}},
  \href{https://doi.org/10.1140/epjc/s10052-017-5157-2}{\emph{Eur. Phys. J. C}
  {\bfseries 77} (2017) 786},
  [\href{https://arxiv.org/abs/1705.07933}{{\ttfamily 1705.07933}}].

\bibitem{GAMBITModelsWorkgroup:2017ilg}
{\scshape GAMBIT Models Workgroup} collaboration, P.~Athron et~al.,
  \emph{{SpecBit, DecayBit and PrecisionBit: GAMBIT modules for computing mass
  spectra, particle decay rates and precision observables}},
  \href{https://doi.org/10.1140/epjc/s10052-017-5390-8}{\emph{Eur. Phys. J. C}
  {\bfseries 78} (2018) 22},
  [\href{https://arxiv.org/abs/1705.07936}{{\ttfamily 1705.07936}}].

\bibitem{Chrzaszcz:2019inj}
M.~Chrzaszcz, M.~Drewes, T.~E. Gonzalo, J.~Harz, S.~Krishnamurthy and
  C.~Weniger, \emph{{A frequentist analysis of three right-handed neutrinos
  with GAMBIT}},
  \href{https://doi.org/10.1140/epjc/s10052-020-8073-9}{\emph{Eur. Phys. J. C}
  {\bfseries 80} (2020) 569},
  [\href{https://arxiv.org/abs/1908.02302}{{\ttfamily 1908.02302}}].

\bibitem{GAMBITCosmologyWorkgroup:2020htv}
{\scshape GAMBIT Cosmology Workgroup} collaboration, J.~J. Renk et~al.,
  \emph{{CosmoBit: A GAMBIT module for computing cosmological observables and
  likelihoods}},
  \href{https://doi.org/10.1088/1475-7516/2021/02/022}{\emph{JCAP} {\bfseries
  02} (2021) 022}, [\href{https://arxiv.org/abs/2009.03286}{{\ttfamily
  2009.03286}}].

\bibitem{Martinez:2017lzg}
{\scshape GAMBIT} collaboration, G.~D. Martinez, J.~McKay, B.~Farmer, P.~Scott,
  E.~Roebber, A.~Putze et~al., \emph{{Comparison of statistical sampling
  methods with ScannerBit, the GAMBIT scanning module}},
  \href{https://doi.org/10.1140/epjc/s10052-017-5274-y}{\emph{Eur. Phys. J. C}
  {\bfseries 77} (2017) 761},
  [\href{https://arxiv.org/abs/1705.07959}{{\ttfamily 1705.07959}}].

\bibitem{GAMBIT:2017snp}
{\scshape GAMBIT} collaboration, P.~Athron et~al., \emph{{Global fits of
  GUT-scale SUSY models with GAMBIT}},
  \href{https://doi.org/10.1140/epjc/s10052-017-5167-0}{\emph{Eur. Phys. J. C}
  {\bfseries 77} (2017) 824},
  [\href{https://arxiv.org/abs/1705.07935}{{\ttfamily 1705.07935}}].

\bibitem{GAMBIT:2017zdo}
{\scshape GAMBIT} collaboration, P.~Athron et~al., \emph{{A global fit of the
  MSSM with GAMBIT}},
  \href{https://doi.org/10.1140/epjc/s10052-017-5196-8}{\emph{Eur. Phys. J. C}
  {\bfseries 77} (2017) 879},
  [\href{https://arxiv.org/abs/1705.07917}{{\ttfamily 1705.07917}}].

\bibitem{GAMBIT:2018gjo}
{\scshape GAMBIT} collaboration, P.~Athron et~al., \emph{{Combined collider
  constraints on neutralinos and charginos}},
  \href{https://doi.org/10.1140/epjc/s10052-019-6837-x}{\emph{Eur. Phys. J. C}
  {\bfseries 79} (2019) 395},
  [\href{https://arxiv.org/abs/1809.02097}{{\ttfamily 1809.02097}}].

\bibitem{GAMBIT:2017gge}
{\scshape GAMBIT} collaboration, P.~Athron et~al., \emph{{Status of the scalar
  singlet dark matter model}},
  \href{https://doi.org/10.1140/epjc/s10052-017-5113-1}{\emph{Eur. Phys. J. C}
  {\bfseries 77} (2017) 568},
  [\href{https://arxiv.org/abs/1705.07931}{{\ttfamily 1705.07931}}].

\bibitem{Athron:2018ipf}
P.~Athron, J.~M. Cornell, F.~Kahlhoefer, J.~Mckay, P.~Scott and S.~Wild,
  \emph{{Impact of vacuum stability, perturbativity and XENON1T on global fits
  of $\mathbb {Z}_2$ and $\mathbb {Z}_3$ scalar singlet dark matter}},
  \href{https://doi.org/10.1140/epjc/s10052-018-6314-y}{\emph{Eur. Phys. J. C}
  {\bfseries 78} (2018) 830},
  [\href{https://arxiv.org/abs/1806.11281}{{\ttfamily 1806.11281}}].

\bibitem{GAMBIT:2021rlp}
{\scshape GAMBIT} collaboration, P.~Athron et~al., \emph{{Thermal WIMPs and the
  Scale of New Physics: Global Fits of Dirac Dark Matter Effective Field
  Theories}},  \href{https://arxiv.org/abs/2106.02056}{{\ttfamily 2106.02056}}.

\bibitem{Hoof:2018ieb}
S.~Hoof, F.~Kahlhoefer, P.~Scott, C.~Weniger and M.~White, \emph{{Axion global
  fits with Peccei-Quinn symmetry breaking before inflation using GAMBIT}},
  \href{https://doi.org/10.1007/JHEP03(2019)191}{\emph{JHEP} {\bfseries 03}
  (2019) 191}, [\href{https://arxiv.org/abs/1810.07192}{{\ttfamily
  1810.07192}}].

\bibitem{Athron:2020maw}
P.~Athron et~al., \emph{{Global fits of axion-like particles to XENON1T and
  astrophysical data}},
  \href{https://doi.org/10.1007/JHEP05(2021)159}{\emph{JHEP} {\bfseries 05}
  (2021) 159}, [\href{https://arxiv.org/abs/2007.05517}{{\ttfamily
  2007.05517}}].

\bibitem{Bhom:2020lmk}
J.~Bhom, M.~Chrzaszcz, F.~Mahmoudi, M.~T. Prim, P.~Scott and M.~White, \emph{{A
  model-independent analysis of $b \to s \mu^{+} \mu^{-}$ transitions with
  GAMBIT's FlavBit}},  \href{https://arxiv.org/abs/2006.03489}{{\ttfamily
  2006.03489}}.

\bibitem{GAMBITCosmologyWorkgroup:2020rmf}
{\scshape GAMBIT Cosmology Workgroup} collaboration, P.~St\"ocker et~al.,
  \emph{{Strengthening the bound on the mass of the lightest neutrino with
  terrestrial and cosmological experiments}},
  \href{https://doi.org/10.1103/PhysRevD.103.123508}{\emph{Phys. Rev. D}
  {\bfseries 103} (2021) 123508},
  [\href{https://arxiv.org/abs/2009.03287}{{\ttfamily 2009.03287}}].

\bibitem{Christensen:2008py}
N.~D. Christensen and C.~Duhr, \emph{{FeynRules - Feynman rules made easy}},
  \href{https://doi.org/10.1016/j.cpc.2009.02.018}{\emph{Comput. Phys. Commun.}
  {\bfseries 180} (2009) 1614--1641},
  [\href{https://arxiv.org/abs/0806.4194}{{\ttfamily 0806.4194}}].

\bibitem{Alloul:2013bka}
A.~Alloul, N.~D. Christensen, C.~Degrande, C.~Duhr and B.~Fuks,
  \emph{{FeynRules 2.0 - A complete toolbox for tree-level phenomenology}},
  \href{https://doi.org/10.1016/j.cpc.2014.04.012}{\emph{Comput. Phys. Commun.}
  {\bfseries 185} (2014) 2250--2300},
  [\href{https://arxiv.org/abs/1310.1921}{{\ttfamily 1310.1921}}].

\bibitem{Staub:2008uz}
F.~Staub, \emph{{SARAH}},  \href{https://arxiv.org/abs/0806.0538}{{\ttfamily
  0806.0538}}.

\bibitem{Staub:2012pb}
F.~Staub, \emph{{SARAH 3.2: Dirac Gauginos, UFO output, and more}},
  \href{https://doi.org/10.1016/j.cpc.2013.02.019}{\emph{Comput. Phys. Commun.}
  {\bfseries 184} (2013) 1792--1809},
  [\href{https://arxiv.org/abs/1207.0906}{{\ttfamily 1207.0906}}].

\bibitem{Staub:2013tta}
F.~Staub, \emph{{SARAH 4 : A tool for (not only SUSY) model builders}},
  \href{https://doi.org/10.1016/j.cpc.2014.02.018}{\emph{Comput. Phys. Commun.}
  {\bfseries 185} (2014) 1773--1790},
  [\href{https://arxiv.org/abs/1309.7223}{{\ttfamily 1309.7223}}].

\bibitem{Semenov:1996es}
A.~V. Semenov, \emph{{LanHEP: A Package for automatic generation of Feynman
  rules in gauge models}},
  \href{https://arxiv.org/abs/hep-ph/9608488}{{\ttfamily hep-ph/9608488}}.

\bibitem{Semenov:2002jw}
A.~V. Semenov, \emph{{LanHEP: A Package for automatic generation of Feynman
  rules in field theory. Version 2.0}},
  \href{https://arxiv.org/abs/hep-ph/0208011}{{\ttfamily hep-ph/0208011}}.

\bibitem{Semenov:2008jy}
A.~Semenov, \emph{{LanHEP: A Package for the automatic generation of Feynman
  rules in field theory. Version 3.0}},
  \href{https://doi.org/10.1016/j.cpc.2008.10.012}{\emph{Comput. Phys. Commun.}
  {\bfseries 180} (2009) 431--454},
  [\href{https://arxiv.org/abs/0805.0555}{{\ttfamily 0805.0555}}].

\bibitem{Mertig:1990an}
R.~Mertig, M.~Bohm and A.~Denner, \emph{{FEYN CALC: Computer algebraic
  calculation of Feynman amplitudes}},
  \href{https://doi.org/10.1016/0010-4655(91)90130-D}{\emph{Comput. Phys.
  Commun.} {\bfseries 64} (1991) 345--359}.

\bibitem{Kublbeck:1990xc}
J.~Kublbeck, M.~Bohm and A.~Denner, \emph{{Feyn Arts: Computer Algebraic
  Generation of Feynman Graphs and Amplitudes}},
  \href{https://doi.org/10.1016/0010-4655(90)90001-H}{\emph{Comput. Phys.
  Commun.} {\bfseries 60} (1990) 165--180}.

\bibitem{Hahn:2000kx}
T.~Hahn, \emph{{Generating Feynman diagrams and amplitudes with FeynArts 3}},
  \href{https://doi.org/10.1016/S0010-4655(01)00290-9}{\emph{Comput. Phys.
  Commun.} {\bfseries 140} (2001) 418--431},
  [\href{https://arxiv.org/abs/hep-ph/0012260}{{\ttfamily hep-ph/0012260}}].

\bibitem{Belyaev:2012qa}
A.~Belyaev, N.~D. Christensen and A.~Pukhov, \emph{{CalcHEP 3.4 for collider
  physics within and beyond the Standard Model}},
  \href{https://doi.org/10.1016/j.cpc.2013.01.014}{\emph{Comput. Phys. Commun.}
  {\bfseries 184} (2013) 1729--1769},
  [\href{https://arxiv.org/abs/1207.6082}{{\ttfamily 1207.6082}}].

\bibitem{Uhlrich:2020ltd}
G.~Uhlrich, F.~Mahmoudi and A.~Arbey, \emph{{MARTY - Modern ARtificial
  Theoretical phYsicist A C++ framework automating symbolic calculations Beyond
  the Standard Model}},
  \href{https://doi.org/10.1016/j.cpc.2021.107928}{\emph{Comput. Phys. Commun.}
  {\bfseries 264} (2021) 107928},
  [\href{https://arxiv.org/abs/2011.02478}{{\ttfamily 2011.02478}}].

\bibitem{Bloor:2021gtp}
S.~Bloor, T.~E. Gonzalo, P.~Scott, C.~Chang, A.~Raklev, J.~E. Camargo-Molina
  et~al., \emph{{The GAMBIT Universal Model Machine: from Lagrangians to
  Likelihoods}},  \href{https://arxiv.org/abs/2107.00030}{{\ttfamily
  2107.00030}}.

\end{thebibliography}\endgroup

\documentclass[a4paper,11pt]{article}
\usepackage{pos}
\usepackage{lmodern}
\usepackage{tikz}
\usepackage{tcolorbox}
\usepackage{amsmath}
\usepackage{slashed}
\usepackage{booktabs}

\title{BSM Developments and Tools}
\author*[a]{Tom\'as E. Gonzalo}

\affiliation[a]{Institute for Theoretical Particle Physics and Cosmology (TTK), RWTH Aachen University, D-52056 Aachen, Germany}

\emailAdd{gonzalo@physik.rwth-aachen.de}

\abstract{In this conference paper I introduce a selection of BSM tools and describe their most recent developments. I choose to focus on tools for the reinterpretation of LHC searches, tools that compute dark matter constraints, beyond-the-Standard-Model (BSM) inference tools, and tools that calculate amplitudes and cross sections directly from the Lagrangian.}

\FullConference{%
 
 The Ninth Annual Conference on Large Hadron Collider Physics - LHCP2021\\
 7-12 June 2021\\
 Online
}

\begin{document}

\begin{flushright}{TTK-21-40}\end{flushright}

\maketitle

\section{Introduction}

We are spoiled for choices when it comes to BSM tools. For many years the community has developed tools for every aspect of BSM physics, from various computations of observables, to abstract level tools able to handle the full Lagrangian of a theory. Due to the large amount of software available, in this conference article I will only describe a selection of tools and their recent developments. I will focus on tools for the recasting of LHC searches, dark matter (DM) tools, statistical inference tools and Lagrangian-level tools.

\section{Reinterpretation of LHC searches}

The reinterpretation of BSM searches has been a profilic field of study in particle physics. Significant advances have been made in the last few years in the interaction between the experimental and theoretical communities, and the exchange of information between them is at its highest, especially after the release of simplified and full likelihoods by CMS and ATLAS, respectively~\cite{LHCReinterpretationForum:2020xtr}. This has had profound positive effects on the whole community as well as the physics outcomes, as it has helped with the long-term preservation of the results and provides motivation for the design of new analyses. Consequently, many recasting tools have been developed, following two different approaches: the \textit{Simplified Models} approach, or the \textit{Fast Simulation} approach.

The \textit{Simplified Models} approach tackles the issue of recasting by decomposing a given model into simplified model topologies and comparing the predicted cross section times branching ratio with that reported by the experiments~\cite{Kraml:2013mwa}. As it does not require any simulation of collisions, this approach is extremely fast and perfectly suitable for large parameter fits. The prominent tool that implements this approach is \textsf{SModelS}, which is in fact responsible for pioneering the idea~\cite{Kraml:2013mwa}. \textsf{SModelS} currently contains the largest database of implemented analyses in the community, with over $100$ run I and II available analyses. With its latest version, \textsf{SModelS} can be interfaced with the full likelihoods provided by the ATLAS experiment~\cite{Alguero:2020grj}. Another notable tool that implements the Simplified Model approach is \textsf{DarkCAST}, a recasting tool for dark photon searches~\cite{Ilten:2018crw}.

On the other hand is the \textit{Fast Simulation} approach, which requires  to perform a Monte-Carlo simulation of the hard scattering process first, then drag the generated events through some form of detector modelling and lastly pass them through an event analysis framework. There are many available event analysis frameworks, some of them with native detector modelling. A non-exhaustive list of these include \textsf{MadAnalysis 5}~\cite{Conte:2012fm, Conte:2014zja, Dumont:2014tja, Conte:2018vmg, Araz:2019otb, Araz:2020lnp}, which can be used with a detector simulation or transfer functions and contains around 40 run I and II analyses; \textsf{CheckMATE}~\cite{Drees:2013wra,Dercks:2016npn}, which contains over 50 run I and II analyses, and was recently extended to support long-lived particle searches in addition to prompt particle searches; \textsf{ColliderBit}~\cite{GAMBIT:2017qxg}, which includes its own detector modelling in the form fast 4-vector smearing (\textsf{Buckfast}), and has a database of around 40 analyses, from runs I and II\footnote{\textsf{ColliderBit} is part of the \textsf{GAMBIT} framework, but it will soon be released as a standalone package}; and \textsf{Rivet}~\cite{Buckley:2010ar, Bierlich:2019rhm}, which has currently ``only'' 30 BSM analyses, but boasts a huge library of SM analyses, upwards of 800. In fact, it has been shown that SM inclusive measurements also set strong constraints on new physics, and the tool \textsf{Contur} has been developed for this purpose~\cite{Butterworth:2016sqg, Amrith:2018yfb, Buckley:2020wzk, Butterworth:2020vnb, Buckley:2021neu}. A comparison of the performance of these tools for a specific CMS search for supersymmetric (SUSY) particles~\cite{CMS:2018kag} was performed~\cite{Brooijmans:2020yij} and strong agreement was seen across all included tools.
  
\section{Dark Matter}

Searches for DM are a strong component of the physics agenda of the LHC experiments, but the strongest constraints on DM models arise often from  direct and indirect detection, as well as from the relic density (RD) of DM, and thus many tools have been developed over the years for that purpose\footnote{A thorough review of many DM tools can be found in \cite{Arina:2020alg}}. In recent times, there is a tendency towards the development of model-independent tools, able to constrain a large variety of BSM models. In addition, many existing and new tools attempt a comprehensive approach to constraints of DM, intending to simultaneously provide constraints from RD, direct and indirect detection. The most employed comprehensive tools developed by the community are \textsf{DarkSUSY 6}~\cite{Gondolo:2004sc, Bringmann:2018lay}, recently expanded to compute the RD beyond the standard kinetic equilibrium; \textsf{MicrOMEGAs}~\cite{Belanger:2001fz,Belanger:2004yn,Belanger:2006is,Belanger:2008sj,Belanger:2010gh,Belanger:2013oya,Belanger:2014vza}, whose most recent upgrade included computations of freeze-in RD; \textsf{MadDM}~\cite{Backovic:2013dpa, Backovic:2015cra, Ambrogi:2018jqj, Arina:2020kko}, part of the \textsf{MadGraph} toolchain; and \textsf{SuperISO Relic}~\cite{Arbey:2009gu, Arbey:2011zz, Arbey:2018msw}, which is widely used for SUSY models.

Even though comprehensive tools can cover all aspects of DM constraints, some more dedicated tools are better suited to compute a single constraint with higher accuracy. In the case of constraints from direct detection (DD), the front runner is \textsf{DDCalc}~\cite{GAMBITDarkMatterWorkgroup:2017fax, GAMBIT:2018eea}, which contains the largest collection of DD experiments to date, and can be used for complete as well as effective DM models, with RG evolution provided by \textsf{DirectDM}~\cite{Bishara:2017nnn, Brod:2017bsw}. In the indirect direction frontier the amount of available tools is too large to recount here. A short list of selected tools include \textsf{Capt'n General}~\cite{Kozar:2021iur}, which imposes constraints on DM capture in the Sun, \textsf{nulike}~\cite{IceCube:2012fvn, IceCube:2016yoy}, which calculates constraints from the DM annihilation to neutrinos, and \textsf{gamLike}~\cite{GAMBITDarkMatterWorkgroup:2017fax}, which deals with DM annihilation to $\gamma$-rays. Lastly one can impose constraints on DM models from cosmological observations. In particular, the energy injected into the Cosmic Microwave Background from DM annihilations in the early Universe can be computed with \textsf{DarkAges}~\cite{Stocker:2018avm}, part of the \textsf{ExoCLASS} branch of \textsf{CLASS}~\cite{Blas:2011aaa}.

\section{Inference and Sampling}

It has long been the goal of many particle theorists to assess the validity of BSM models in light of data. However, the vast quantities of data produced by the experiments, and the increasingly large parameter spaces of the models, makes this a difficult task. The common practices of overlaying exclusing regions and sampling the parameter space using random or grid samplers are insufficient and much information is lost along the way. Composite likelihoods that keep all the information about experimental data, smart sampling strategies to explore large and intricate parameter spaces~\cite{DarkMachinesHighDimensionalSamplingGroup:2021wkt}, and a rigorous treatment of statistics are critical features that significantly improve and simplify this task~\cite{AbdusSalam:2020rdj}. And for this purpose many global fitting tools have been developed that include those features and that are targeted, though not exclusively, to BSM models. 

Bayesian inference has been a staple in astrophysics for many years, and there are many public tools developed for that purpose, but in the field of BSM phenomenology the tendency has been to develop private fitters for specific purposes, e.g. flavour fits, neutrino oscillations, SUSY, etc. Only in recent years various public inference tools have been released. The first ``wave'' of publicly available global fitters, such as \textsf{Fittino}~\cite{Bechtle:2012zk}, \textsf{SuperBayes}~\cite{RuizdeAustri:2006iwb} or \textsf{EasyScanHEP}~\cite{Han:2016gvr}, were developed to fit various SUSY models to LHC data, and some are still in use today.

The revolution in the field of BSM inference came with the advent of model-independent tools. The front-runner for SM fits is \textsf{HEPFit}~\cite{DeBlas:2019ehy}, built upon a Bayesian Markov chain Monte Carlo, which includes a vast library of electroweak precision, Higgs and flavour observables, and can also work with BSM models~\cite{Chiang:2018cgb,Chowdhury:2017aav}. Nevertheless, the indisputable leader of BSM inference studies is the \textsf{GAMBIT} tool~\cite{GAMBIT:2017yxo,Kvellestad:2019vxm}. \textsf{GAMBIT} contains a huge library of likelihoods and observables for collider physics~\cite{GAMBIT:2017qxg}, dark matter~\cite{GAMBITDarkMatterWorkgroup:2017fax}, flavour~\cite{GAMBITFlavourWorkgroup:2017dbx}, precision observables~\cite{GAMBITModelsWorkgroup:2017ilg}, neutrino physics~\cite{Chrzaszcz:2019inj} and cosmology~\cite{GAMBITCosmologyWorkgroup:2020htv}, as well as an extensive database of models and sampling strategies, both frequentist and Bayesian~\cite{Martinez:2017lzg}. \textsf{GAMBIT} has been employed in studies of a large variety of BSM modes, including SUSY models~\cite{GAMBIT:2017snp,GAMBIT:2017zdo,GAMBIT:2018gjo}, Higgs-portal DM models~\cite{GAMBIT:2017gge,Athron:2018ipf,GAMBIT:2018eea}, EFT DM models~\cite{GAMBIT:2021rlp}, axion and axion-like particle models~\cite{Hoof:2018ieb,Athron:2020maw}, models with right-handed neutrinos~\cite{Chrzaszcz:2019inj}, fits of neutral flavour anomalies~\cite{Bhom:2020lmk} and cosmological fits of neutrino masses~\cite{GAMBITCosmologyWorkgroup:2020rmf}.

\section{Lagrangians and Model Building}

In addition to the software developed for the computation of observables, there is a small collection of tools that go one step further, and attempt to bridge the gap between the full Lagrangian of the BSM theory and the required amplitudes and cross sections. Some of these tools, such as \textsf{FeynRules}~\cite{Christensen:2008py, Alloul:2013bka}, \textsf{SARAH}~\cite{Staub:2008uz,Staub:2012pb,Staub:2013tta} and \textsf{LanHEP}~\cite{Semenov:1996es,Semenov:2002jw,Semenov:2008jy}, compute the Feynman rules directly from the Lagrangian, while others, such as \textsf{FeynArts/FormCalc}~\cite{Mertig:1990an, Kublbeck:1990xc, Hahn:2000kx} or \textsf{CalcHEP}~\cite{Belyaev:2012qa}, use those to compute the required amplitudes. Although widely used, the main disadvantage of most of these tools is that they are written in \textsf{Mathematica}, which is a proprietary software framework and it is notably slower than other interpreted or compiled languages. Very recently a new tool has been released, \textsf{MARTY}~\cite{Uhlrich:2020ltd}, a fully open-source framework that uses symbolic computations in \textsf{C++} to obtain amplitudes and cross sections from the Lagrangian for any BSM theory. Finally, the last step in the BSM toolchain is to automatically connect the amplitudes, cross sections, etc, obtained by the Lagrangian-level tools, with the calculation of likelihoods and observables. This step is most often done by hand, by manually embedding the code or expressions outputted by the above tools where needed. Fortunately, an automatic integration of this step can now be achieved with \textsf{GUM}~\cite{Bloor:2021gtp}. By interfacing with \textsf{SARAH} and \textsf{FeynRules}, \textsf{GUM} extracts all information about particles, parameters and expressions, auto-generates code in \textsf{GAMBIT}, and uses it to impose DM, collider and vacuum stability constraints. Using both \textsf{GUM} and \textsf{GAMBIT} one can start with a Lagrangian and end with a complete likelihood analysis without writing a single line of code.

\section{Summary}

In this conference article I have introduced a selection of tools for BSM physics and described their most recent developments. The state of the communication between the theory and experimental communities is currently at its best, and that improves the quality and relevance of the results from recasting tools. There is a community trend towards the development of model-independent tools, as it is the case for some of the DM comprehensive tools, as well as the leading global fitting frameworks, \textsf{HEPFit} and \textsf{GAMBIT}. And lastly, with \textsf{MARTY} and \textsf{GUM} we have managed to complete the BSM toolchain, from Lagrangian to likelihoods, using only open-source software.

\bibliographystyle{JHEP}
\bibliography{LHCPGonzalo}

\end{document}